Given demand is P=2000-4Q & supply is P=1000+Q. The coinsurance rate is 0.25. what is the equilibrium quantity and price, consumer/producer surplus and deadweight loss of moral hazard( graph would help)
1
Expert's answer
2021-05-08T14:41:27-0400

The equilibrium quantity and price are:

2000 - 4Q = 1000 + Q,

5Q = 1000,

Q = 200 units,

P = 1000 + 200 = 1200.

Consumer surplus is:

CS = 0.5×(2000 - 1200)×200 = 80,000.

Producer surplus is:

PS = 0.5×(1200 - 1000)×200 = 20,000.

With the coinsurance the deadweight loss will occur.